[Brahms-l] Shift report 20050206 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Feb 06 2005 - 16:02:59 EST
1. Good store when I came, dump scheduled for 12:00. After dumped the beam, we got another good store around 1:00pm, dump scheduled for 5:50pm

2. Runs:

FS@12deg@A430, MRS@45deg@B2255 
RunNo   Trig6     Trig3
12831 - 211k      107k 
12832 - 152k      119k
12833 - 175k      153k

FS@4deg@B1219  MRS@45deg@B2255
RunNo   Trig2      Trig3
12834 -  OFF      84k (maybe not good, because the magnet people turned off the D5 accidently when they came to change the polarity of FS magnet) 
12836 -  42k      OFF (D5 got problem to ramping up to 2255 after been turned down accidently.)
12840 -  99k      43k
12841 - RunServer problem
12842 -  228k     81k 
12843 -  133k     75k
12844 -  ongoing

TPC pedestal run     12838
non-TPC pedestal run 12839

3. T3 had problem, see the previous shift report. And I did not see the bubbles in T1 and T2. See e-log for detail info.



_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 6 16:03:20 2005

This archive was generated by hypermail 2.1.8 : Sun Feb 06 2005 - 16:03:38 EST